use elimination to solve the system of
7x+23y=12
3x+23y=-8

Answer by checkley77(12844)

7x+23y=12
-(3x+23y=-8) SUBTRACT
----------------
4X=20
X=20/4
X=5 ANS.
7*5+23Y=12
35+23Y=12
23Y=12-35
23Y=-23
Y=-23/23
Y=-1 ANS.
PROOF:
3*5+23*-1=-8
15-23=-8
-8=-8